OpenStudy (anonymous):

@sammixboo What is the approximate volume of the cylinder? height = 6.4 m; diameter = 4.6 m

sammixboo (sammixboo):

Do you know the formula to finding the volume of a cylinder?

OpenStudy (anonymous):

@horsegirl27 @sammixboo is it d and height?

sammixboo (sammixboo):

It is \(\rm \color{blue}{V = \pi r^2h}\) Where \(\rm \color{blue}{\pi}\) = 3,14 \(\rm \color{blue}{r}\) = radius = half of diameter \(\rm \color{blue}{h}\) = height First of all, what is half of the measurement of the diameter?

sammixboo (sammixboo):

4.6/2 = ?

OpenStudy (anonymous):

2.3

sammixboo (sammixboo):

Correct, so let's plug in the values \(\rm \color{blue}{V = 3.14 \times 2.3^2\times 6.4}\) First, we must do the exponent \(\color{blue}{2.3^2}\) \(\rm \color{blue}{2.3^2}\) \(\rightarrow\) \(\rm \color{blue}{2.3\times2.3}\)

sammixboo (sammixboo):

So what is 2.3 x 2.3?

OpenStudy (anonymous):

5.29

sammixboo (sammixboo):

Correct, so \(\rm \color{blue}{V = 3.14 \times 5.29\times 6.4}\) Now multiply. Starting with 3.14 x 5.29